OpenStudy (anonymous):

integrate the following:

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ \sqrt{1-sinx} }dx\]

OpenStudy (experimentx):

1 - sin(x) = (sin(x/2) - cos(x/2))^2

OpenStudy (experimentx):

sin(x/2) - cos(x/2) = sqrt(2)sin(3pi/4 + x)

OpenStudy (experimentx):

or make use of Weierstrass substitution